? ? œ ?? L H N = HL Ω Ω ⊂

Z^{2} Ω pixel

k ∈ [[1; N]] x_{k} =

(i, j)_{k} ∈

Ω v_{k} =

v(i, j)_{k}

1 N v̄ = (v_{k})_{k∈[[1;N]]}

Ω ū ū

ˆū = (û_{k})_{k∈[[1;N]]}

ū v̄ v̄ = ū+ σϵ ϵ ∈

R^{N} ū σ v̄ ϵ ū p(v|u)

photosites ? ? ? ? additif

σ^{2} σ > 0

p(v|u) = \frac{1}{\sqrt{2}πσ} e^{− \frac{(v−u)^{2}}{2σ^{2}}}

multiplicatif L

p(v | u) = \frac{L^{2}v^{(L−1)}e^{−L \frac{v}{u}}}{Γ(L)u^{L}}

E [v] = u σ^{2} =

\frac{u^{2}}{L}

L L = 1 u^{2} v sel et poivre impulsionnel P

p(v | u) = { P 2+(1−P)si

et \frac{P}{2} +

(1− P)siet \frac{P}{2} siet

(≠ 0) / (\overline{2si}et)

≠ 255 v = u u ∉ {0, 255}

bruit de grenaille

p(v | u) = e\frac{u^{v}}{v!}

speckle Poisson ?? ?

P SNR = 10log_{10} ( \frac{D^{2}}{\frac{1}{N} ∑_{k<N} (v_{k} − u_{k})^{2}} )

v̄ ū D et al. ? [0; 1]

gilles/T HESE/codes/ny256.png[Bruitgaussienσ =

25 gilles/T HESE/codes/ny256_{g}auss25.png[Bruitimpulsionnel25%, P SNR =